There are approximately 4lbs of muscle for every 7lbs of body weight of a certain animal. For a 175 lb. animal, how much of the weight is muscle?
Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

There are approximately 4lbs of muscle for every 7lbs of body weight of a certain animal. For a 175 lb. animal, how much of the weight is muscle?
----------------------
x/175 = 4/7
---
x = 175(4/7)
---
x = 100 lbs of muscle
========================
Cheers,
Stan H.

Answer by rfer(16322)

4/7=x/175
7x=700
x=100 lbs